OpenStudy (qwertykingghlpo):

Use the quadratic formula to solve the equation. -5x^2 - 14x + 3 = 0 Select the correct choice below and, if necessary, fill in the answer box to complete your choice. A. x =? B. The solution is not a real number.

MsBrains (ms-brains):

Any ideas? x3

OpenStudy (qwertykingghlpo):

not really :/

MsBrains (ms-brains):

Is this all of the question? A is x=?

OpenStudy (qwertykingghlpo):

yes

OpenStudy (qwertykingghlpo):

im unsure?...what happened?

OpenStudy (mjdennis):

Hey, @Ms-Brains, note that he was instructedto use the quadratic formula here.

OpenStudy (mathmale):

qwertyking: I'd suggest you look up and copy down the quadratic formula. You will likely need this formula again and again. Next, take a look at the given polynomial (a quadratic): -5x^2 - 14x + 3 = 0. It has three coefficients, which here happen to be -5, -14 and 3. We label these a=-5, b=-14, and c=3. Substitute these values into the quadratic formual (which you have already looked up and copied down for reference). simplify your results. x=?? (there must be TWO roots / solutions / zeros).

OpenStudy (qwertykingghlpo):

what do you mean @mjdennis ?

OpenStudy (qwertykingghlpo):

I know the quadratic formula already.

OpenStudy (mathmale):

Great. type it in here, please. x= ?

OpenStudy (qwertykingghlpo):

x = -b + or 1 the square root of the discrimant (b^2 - 4ac)....all of this over 2a

OpenStudy (qwertykingghlpo):

I mean + or -

OpenStudy (mathmale):

x = -b plus or minus the square root of the discrimant (b^2 - 4ac)....all of this over 2a Good. Now, b=-14, so, after substitution, you get: x = -[-14] plus or minus the square root of the discriminant ( [-14]^2 - 4ac)....all of this over 2a. We still have to insert the known values of a and c. But first, simplify this: x = -[-14] plus or minus the square root of the discriminant ( [-14]^2 - 4ac)....all of this over 2a.

OpenStudy (qwertykingghlpo):

okay so -14 squared is -196. So we are looking for -196 - 4ac over 2a....

OpenStudy (mathmale):

Not quite. -14 squared is (-14)^2 = ? Those parentheses are mandatory. What about the "-b" part of the quadratic equation? Please be certain to include it.

OpenStudy (qwertykingghlpo):

woops postive 196

OpenStudy (mathmale):

Yes, and so \[x=\frac{ -(-14)\pm \sqrt{196-4(a)(c)} }{ 2(-5) }\]

OpenStudy (mathmale):

What next? You must insert the proper values of a and c.

OpenStudy (qwertykingghlpo):

now how do I find that?

OpenStudy (mathmale):

a=-5, so merely substitute (-5) for a; c=3, so subst. (3) for c.

OpenStudy (qwertykingghlpo):

how did you know?

OpenStudy (mathmale):

You must follow the quadratic formula as tho' it were a road map. The values of a, b and c come from the given quadratic equation, y=-5x^2-14x+3.

OpenStudy (mathmale):

As before, a=-5, b=-14, c=3.

OpenStudy (mathmale):

Substituting c=3, we get:\[x=\frac{ -(-14)\pm \sqrt{196-4(a)(3)} }{ 2(-5) }\]

OpenStudy (mathmale):

Finally, where you see (a), substitute (-5).

OpenStudy (mathmale):

Please share everything you do from this point on.

OpenStudy (qwertykingghlpo):

what about the a next to the 4?

OpenStudy (mathmale):

that's the "a" I was talking about. Replace that (a) with (-5) now, please.

OpenStudy (qwertykingghlpo):